OpenStudy (anonymous):

A cupboard contains 5 pairs of shoes.4 shoes are drawn one by one at random. Find the probability that their atleast one pair of shoes is drawn?

OpenStudy (science0229):

How many shoes are there in total?

OpenStudy (anonymous):

5 pairs of shoes

OpenStudy (science0229):

There are 2 shoes per pair. So there are 10 shoes in total, right?

OpenStudy (anonymous):

yes

OpenStudy (science0229):

Have you learned combinations, yet?

OpenStudy (anonymous):

yes i know combination

OpenStudy (science0229):

Then you can easily find the number of ways to choose 4 shoes out of 10 shoes.

OpenStudy (anonymous):

it is not directly 4 out of 10 shoes. we need to choose atleast one pair, the answer should be 13/21. can u tell the answer is coming

OpenStudy (anonymous):

can u tell how the anwer is coming

OpenStudy (science0229):

We're getting to the answer. The number of ways to choose 4 shoes out of 10 shoes is in the denominator.

OpenStudy (anonymous):

yes what is the numerator

OpenStudy (science0229):

Now, it's saying to find the probability that you pick at least 1 pair of shoe.

OpenStudy (science0229):

If you do 1-(probability that no pair of shoe come out), then that's the same thing as the probability to pick at least 1 pair of shoe.

OpenStudy (anonymous):

how to do that

OpenStudy (science0229):

Say that the shoes are a1,a2,b1,b2,c1,c2,d1,d2,e1,e2,f1,f2

OpenStudy (anonymous):

ok ok sir

OpenStudy (science0229):

Wait. Sorry. say that the shoes are a1,a2,b1,b2,c1,c2,d1,d2,e1,e2

OpenStudy (science0229):

we have to remember that the order doesn't matter.

OpenStudy (science0229):

When we pick the first shoe, it can be any of the 10 shoes, right?

OpenStudy (anonymous):

ok ok

OpenStudy (science0229):

When we pick the second shoe, we already picked 1 and we can't pick the pair of it, so it has to be from 8 shoes, right?

OpenStudy (anonymous):

yes

OpenStudy (science0229):

using similar way, the 3rd time is 6 and the 4th time is 4.

OpenStudy (science0229):

So you might think that there are 10*8*6*4 ways to pick 4 shoes without any pairs.

OpenStudy (anonymous):

yes that's right

OpenStudy (science0229):

BUT, I mentioned that order doesn't matter.

OpenStudy (science0229):

For example, (a1,b1,c1,d1) and (b1,a1,c1,d1) are considered the same.

OpenStudy (anonymous):

ok

OpenStudy (science0229):

So for every combination we counted 4! times. So the actual ways to count without picking any pairs can be calculated as (10*8*6*4)/4!

OpenStudy (science0229):

Following me so far?

OpenStudy (anonymous):

yes

OpenStudy (science0229):

So, the probability to choose 4 shoes without any pairs is\[\frac{ \frac{ 10*8*6*4 }{ 4! } }{ 10C4 }\]. 10C4 is 10 combination 4

OpenStudy (science0229):

Calculate that to get 8/21. So the probability to pick 4 shoes with at least 1 pair is 1-(8/21)=13/21

OpenStudy (anonymous):

i understood everything but why there should be division of 4 factorial. can u explain

OpenStudy (science0229):

Take these 4 combinations a1,b1,c1,d1 b1,a1,c1,d1 a1,b1,d1,c1 c1,a1,b1,d1

OpenStudy (science0229):

Because the order doesn't matter all 4 of them are same, right?

OpenStudy (anonymous):

thank u very much sir

OpenStudy (science0229):

So for that, the number of ways to order those 4 shoes are 4!. Because we counted 4! combinations as same, we have to divide by 4!

OpenStudy (science0229):

you're welcome :)